# A local, many-worlds, model of quantum correlations with finite information flow

### v2

@@ I Introduction
A pragmatic approach is to assume that our expectation of being in a particular branch is given Born’s rule, which is adopted as a postulate. A concrete derivation of the rule from a branch counting is obtained by assuming that there is a multitude of branches such that the ratio between the branches in |+ket|$\rangle$| + ⟩ and those in |ket|$\rangle$| - ⟩ is about $|\psi_{+}/\psi_{-}|^{2}$, as proposed by Saunders saunders . Since this ratio is a real number, the Born rule is recovered exactly in the limit of infinite branches. Thus, if a system splits into a finite number of macroscopically distinct states, each one actually corresponds to an infinite multitude of parallel worlds. Hereafter, in the context of many-worlds theory, we just assume that branches are weighted by an amplitude and, for the sake of concreteness, we also refer to the branch counting of Ref. saunders .
While (single-world) ontological theories are probabilistic, the many-worlds theory is $\psi$-ontic and deterministic. Randomness arises from the subjective perspective of an observer following a single branch of the tree. In this paper, we explore how integrating the randomness of ontological theories with many-worlds branching can yield a more economical framework while preserving an advantage of Deutsch-Hayden (DH) many-worlds theory deutsch – its claimed ability to provide a local, realistic picture of the world(s). The argument supporting this ability is clearly illustrated by a simple local realistic model of a ‘nonlocal’ Popescu-Rohrlich (PR) box brassard . The core idea of the argument is to delay the buildup of correlations until the parties compare their outcomes at a meeting point. However, DH theory requires branch weights and an infinite information flow to the meeting point deutsch . Employing the branch counting of Ref. saunders , also the number of branches turns out to be infinite.
In this paper, by adopting a hybrid framework that incorporates both randomness and branching, we show that just one classical bit of information about the chosen measurement and two “equally weighted” branches are sufficient to simulate local projective measurements on maximally entangled qubits. Randomness allows us to derive the quantum probabilities more naturally and economically from a simple unweighted count of branches and an ensemble average. Furthermore, the $\psi$-epistemic nature of the model leads to a drastic reduction of the information flow. While in the many-worlds theory quantum states are part of the ontology, which must therefore contain infinite classical information, in our model the quantum state is rather encoded in the statistics of the ontic states.
Our model directly builds on the Toner-Bacon model, which simulates maximally entangled states using one bit of communication and shared randomness toner . By selecting a branch at random, the model reproduces the same statistics as the single-world model in Ref. gisin , in which outcomes are generated through shared randomness and the single use of a nonlocal PR box.
II Nonlocal correlations in many-worlds theory
@@
In Ref. deutsch , Deutsch and Hayden proposed a many-worlds, local description of nonlocal quantum correlations using the Heisenberg picture. Here, we present a highly simplified caricature of their argument. While this oversimplification may make the argument seem somewhat trivial, it nevertheless captures its essence.
@@
The qubits are sent to two spatially separated parties, Alice and Bob. Each party independently chooses to perform a projective measurement, represented by the Bloch vectors $\vec{a}$ and $\vec{b}$, respectively. The outcomes, ±$\pm 1$, of each measurement occur with equal probability. In the DH framework, this means that qubits locally split into two branches with equal weights. Rephrasing in Saunders’ terms, there is a multitude of realities in half of which a party observes one of the two outcomes. On each side, this process occurs without the need to know what is happening on the other side. After the measurements, the parties decide to compare their results. To do that, they have to meet at the same place or send some communication through a classical channel. After multiple running of the experiment over different entangled states, they find out that the outcomes, $s_{A}\in\{-1,1\}$ and $a_{B}\in\{-1,1\}$, are distributed according to the probability distribution
$P(s_{A},s_{B}|\vec{a},\vec{b})=\frac{1}{2}\left(1-s_{A}s_{B}\vec{a}\cdot\vec{b}\right),$ (2)
which violate the Bell inequalities. While this violation leads to a break of local realism in a single-world scenario, the DH framework evades this conclusion as follows. When Alice meets with Bob, four branches emerge such that each one is weighted with an amplitude whose modulus square is exactly $P(s_{A},s_{B}|\vec{a},\vec{b})$. Crucially, these weights are not in conflict with the previous weight assignments. Rephrasing in Saunders’ terms, the multitudes of realities of each party are paired in such a way that the fraction of realities experiencing the outcomes $s_{A}$ and $s_{B}$ is equal to $P(s_{A},s_{B}|\vec{a},\vec{b})$. This argument is illustrated in Ref. brassard for nonlocal PR boxes and further discussed in the general quantum case in Ref. charles . The existence of many local realities allows us to delay the buildup of the correlations until the parties meet or communicate each other.
It is fundamental to remark that, in the DH theory, each party has to carry the information about the performed measurement at the meeting point in order to get the right pairing. This information is clearly infinite, which is a consequence of the $\psi$-ontic nature of the theory.
III A model with finite information flow

### v3

@@ III.1 Multiple instances of a macroscopic quantity
$\lambda=(s_{1},s_{2},\dots,s_{m},\eta),$ (2)
[…]
Since the preparation of the quantum state does not uniquely determine the ontic state $\lambda$, the $m$-tuple of outcomes $(s_{1},s_{2},\dots,s_{m})$ generally emerges from the measurement with probability $\rho(s_{1},s_{2},\dots,s_{m})$. Like in many-worlds theory, a single observer will experience one of the $m$ outcomes. In a single-world theory, one would have a distribution $\rho(s)$ over a single outcome $s$, consistent with Born’s rule. In our $m$-instances scenario, however, we must introduce an additional assumption to connect the joint distribution $\rho(s_{1},s_{2},\dots,s_{m})$ with the empirical data observed by a single observer:
Assumption (Equiprobability of instances). The observer is equally likely to be in either reality, without any weight attached to the instances.
This is motivated by the principle of indifference (also known as the principle of insufficient reason). The same assumption has been made in Refs. hall ; brassard . Accordingly, the probability of observing outcome $s$ is
$\rho(s)=\frac{1}{m}\sum_{k}\sum_{s_{1},\dots,s_{m}}\delta_{s_{k},s}\rho(s_{1},\dots,s_{m})$ (3)
This is equivalent to a simple unweighted counting rule: we count how many of the $m$ realities yield outcome $s$ (from 0 to $m$ realities), average over many runs, and divide by $m$. This evaluated $\rho(s)$ must agree with Born’s rule.
Two points are crucial here. First, the number of instances with given outcome $s$ fluctuates from run to run, such that the observed frequencies converge to those predicted by quantum theory as the number of runs goes to infinity. If the theory were deterministic, one would observe only frequencies equal to a rational number in $\{0,1/m,2/m,\dots,1\}$ (the latter with $s=s_{1}=s_{2}=s_{3}=\dots=s_{m}$). Second, in each individual run, not all outcomes need to be realized, in contrast to the many-worlds framework, where all the branches of the wave function are instantiated. In our example, at most $m$ distinct outcomes can occur in a single run (only one if $s_{1}=s_{2}=\dots=s_{m}$). If $m<n$, where $n$ is the number of branches of the wave function, we would have at most $m$ distinct outcomes per run, not the $n$ outcomes realized in the many-worlds theory. Even if $m>n$, the number of distinct realities may be smaller than $n$ in some runs of the experiment. In particular, if there exists a branch of the wave function with an extremely small probability $p$, this means that, based on our knowledge of the system, there is an extremely small probability (not greater than $mp$) that some instance realizes that branch. Thus, it is not guaranteed that such an instance exists in every run of the experiment. This is trivially true in a single-world ontological theory with only one existing reality. This departs sharply from the deterministic realization of all the branches in many-worlds theories. Provided that the ontic state $\lambda$ does not contain $\psi$, the framework also departs from dBB theory, where there are branches which are physically real, but empty of observers.
It is trivial to prove that this framework is consistent with quantum mechanics. In the de Broglie-Bohm theory with one instance ($m=1$), $\eta$ is identified with the wave function and $s_{1}$ represents the position of all the particles in the laboratory (including the particles of the pointer). A naive many-instances theory can be trivially built by introducing $m$ trajectories in the pilot-wave theory such that the position of the particles in each instance is distributed according to Born’s rule. Whereas such many-instances extensions of the de Broglie-Bohm theory shows that the framework does not conflict with quantum mechanics, they have no appealing motivation. The dBB theory consistently solves the measurement problem (which was the motivation of the theory), but raises new problems (nonlocality bell ) and critics (existence of “empty” branches brown ). Introducing additional trajectories does not weaken these new concerns. It is decisively more interesting to investigate $\psi$-epistemic theories of multiple instances, in which $\eta$ is not identified with the quantum state or, more generally, $\lambda$ does not completely define the quantum state.
At this point, it may be useful to clarify the terms instances, realities and branches used in our context. This will further highlight the similarities between our general framework and many-worlds theory – the latter being a limiting case of the former. Later, we will discuss how to combine different macroscopic quantities. This will allow us to introduce the concepts of separability and locality.
III.2 Branches versus instances
Branches, worlds, and realities are interchangeable terms in the context of Everett’s many-worlds theory. When the wave function goes into a superposition of macroscopically distinct states, we say that one world branches into separate worlds or macroscopically distinct realities. Thus, by definition, two branches, worlds or realities differ by some physical quantity. For example, if the physical quantity has $n$ different values and each value has a nonzero probability amplitude, we say that we have $n$ different branches or worlds, although they may have different weights.
In our context, there is a distinction between realities, or instances, and the branches identified by the values taken by physical quantities. Instances are meant as realities with their own private ontic variables $s_{k}$, but these variables may take identical values in different instances. Consequently, we say that the system branches into two distinguishable realities when some physical quantity in different instances evolves from one shared initial value to two distinct values. Let us further clarify this point by discussing an alternative representation of the ontological theory, which aligns more closely with the narrative of the many-worlds theory.
[…]
$\lambda=(n_{1},\dots,n_{M},\eta),$ (4)
[…]
The new representation using occupation numbers is more akin to the usual picture of a many-worlds theory. The label $i$ identifies a branch, whereas the occupation numbers, which are integers, replace the quantum probabilities weighting the branches of many-worlds theory. Provided that the number of branches exceeds the total occupation number, it is clear that many branches have zero occupation number (zero instances). In a single-world ontological theory, the total occupation number is just one, so that only one branch is realized.
Our weighting of branches with an occupation number also somehow resembles Saunders’ branch counting saunders . Saunders argues that the quantum states continuously undergo a huge number of branching at the microscopic level. These microscopically distinguishable states result into few indistinguishable macroscopic states. By partitioning macroscopic branches into suitable sets of microscopic branches with equal weight, Born’s rule is recovered through a simple count of microscopic branches in the limit of infinite partitioning. Saunders’ microstates and macrostates are akin to instances and branches in our terminology. As in our case, each value of a macroscopic observable is weighted by an integer. The mark difference is that Saunders’ weights must go to infinity to recover Born’s rule because of determinism. Conversely, the weights $n_{i}$, which are stochastic variables, are not proportional to the quantum probabilities in a single run – Born’s rule is recovered only on average. This is what happens in every single-world ontological theory, in which only one value has occupation number equal to $1$, whereas the others are equal to zero. The value taken by the single instance changes stochastically from run to run. Consider, for example, the de Broglie-Bohm theory, in which the particle positions are probabilistically distributed according to Born’s rule.
III.2.1 Many-worlds theory in the limit $m\rightarrow\infty$
[…]
$\bar{n}_{i}=|\psi_{i}|^{2}$ (5)
in each run. The quantity $\bar{n}_{i}$ is the fraction of instances in the branch $i$, which resembles Saunders’ branch counting. This quantity offers a convenient way to explain why branches with greater weights are more likely to be experienced – they are associated with a higher multitude of instances, so that the previously stated assumption of equiprobability of instances holds. In the limit $N_{t}\rightarrow\infty$, the framework with constraint (5) is de facto equivalent to many-worlds theory with the quantities $\bar{n}_{i}$ as extra structure. The constraint (5) may be justified by the diverse derivations of Born’s rule given in many-worlds theory. For example, the constraint can be derived by assuming that $\bar{n}_{i}$ is a function of $|\psi_{i}|$ and that an additivity property holds, along the same lines as Everett everett . Namely, the additivity property states that the probability weight of $\sum_{k}a_{k}|\phi_{k}\rangle$ is equal to the sum of the probability weights of $a_{1},a_{2},\dots$, where $|\phi_{1}\rangle,|\phi_{2}\rangle,\dots$ are orthonormal vectors. This property guarantees consistency of the constraint under unitary evolutions – namely, it ensures the normalization of the probabilities over time.
It can be interesting to investigate whether constraint (5) can be imposed only on average by retaining the stochasticity of the framework in the limit $N_{t}\rightarrow\infty$.
III.2.2 Conservation of the number of instances
The representation using occupation numbers might raise the question of whether the total number of instances can be created or annihilated, analogous to particles in quantum field theory. Given the analogy between occupation numbers and the quantum probabilities weighting the branches, the conservation of the total number of instances appears as a natural hypothesis, mirroring the conservation of the norm of the wave function. Indeed, the creation of new instances would correspond to the ex novo creation of macroscopic systems – such as observers and measuring devices. This would be peculiar, especially considering that each instance is regarded as a physically existing system. Such creation would imply an increase in energy. Indeed, branching in the many-worlds theory does not involve the creation of new systems, but rather a divergence into different states of the same system. Similarly, in our framework, branching corresponds to a fixed number of instances whose states diverge from a shared initial state.
III.3 Combination of two macroscopic quantities
Before introducing separability and locality in the theory, let us discuss the rule for combining two macroscopic quantities, say ${\cal M}_{A}$ and ${\cal M}_{B}$, generated by performing two different measurements on the quantum system $\cal Q$}. Each of these quantities are, for example, the pointer of some measuring device. There are different ways of generating ${\cal M}_{A}$ and ${\cal M}_{B}$. Let us analyze three cases:
  1. 1.
    One observer, say Alice, performs first one measurement generating the macroscopic quantity ${\cal M}_{A}$, then she performs a second measurement generating ${\cal M}_{B}$.
  2. 2.
    Alice performs a measurement, generating ${\cal M}_{A}$, then departs from the measured system. Subsequently, another party, Bob, performs a second measurement, generating ${\cal M}_{B}$. Alice and Bob have not interacted with each other in the interval between the measurements.
  3. 3.
    Given two entangled systems ${\cal Q}_{A}$ and ${\cal Q}_{B}$ that are spacelike-separated, Alice and Bob perform a local measurement on ${\cal Q}_{A}$ and ${\cal Q}_{B}$, respectively. Eventually, they meet to compare their results.
Let us assume that there are $m$ realities of Alice. In the first case, when Alice performs the first measurement, the pointer of one measuring device in the $k$-th instance moves to the position $s_{A,k}$, which is observed by Alice. After a second measurement is performed, the pointer of another device moves to position $s_{B,k}$, which is again observed. Since the number of realities of Alice is preserved, in the end, the overall quantity ${\cal M}={\cal M}_{A}\oplus{\cal M}_{B}$ has $m$ instances with value $(s_{A,k},s_{B,k})$ in the $k$-th instance. Thus, the value $s_{k}$ in Eq. (2) is identified with the pair $(s_{A,k},s_{B,k})$. It is worth to note that the combination of the quantities corresponds to a direct sum, rather than a tensor product.
The second case is more intricate. Let us assume that Bob has a number $m^{\prime}$ of instances. Since Alice and Bob never interact in the interval between the measurements, all we can say immediately after the last measurement is that Alice has $m$ instances of $s_{A}$ and Bob $m^{\prime}$ instances of $s_{B}$. There is no reason for pairing their respective outcomes as long as they stay far away. The ontic state is
$\lambda=(s_{A,1},s_{A,2},\dots,s_{A,m},s_{B,1},s_{B,2},\dots,s_{B,m^{\prime}},\eta),$ (6)
in which the ordering of $s_{A}$ and $s_{B}$ establishes no pairing rule of Alice and Bob’s instances. In this respect, Peres’ dictum ”unperformed measurements have no result” peres holds. Only when Alice and Bob meet, a pairing of their instances is defined. Since the number of instances of each party is preserved and one party eventually experiences one definite macroscopic state of the other party, the pairing must be one-to-one, so that $m^{\prime}=m$. For example, if Alice had fewer instances, some of them would need to pair with more than one of Bob’s. Such a mismatch would generally require the creation of Alice’s instances.
The specific way in which the pairing occurs depends on the ontic state at the moment the meeting takes place. The pairing is understood as a physical process in which one instance of Alice interacts with only one instance of Bob. Note that if the theory is deterministic, the ontic state contains information about the pairing even prior to the meeting of the parties. It is akin to saying that a neutron and a proton will bind if they are brought sufficiently close together. We seek a local mechanism for the pairing, which, however, may be mathematically defined in a nonlocal way at any given time.
The framework may be generalized to infinite instances. In this generalization, it is required that the cardinality of the set of instances is identical for each quantity (e.g. countable, continuum and so on).
Summarizing, we have the following properties
  • Each macroscopic quantity has a fixed number $m$ of instances (more generally, the set of instances has fixed cardinality).
  • If two quantities ${\cal Q}_{A}$ and ${\cal Q}_{B}$ are jointly known, then there is a one-to-one correspondence between the instances of ${\cal Q}_{A}$ and the instances of ${\cal Q}_{B}$. This correspondence defines a combination ${\cal Q}_{A}\oplus{\cal Q}_{B}$ with $m$ instances. The correspondence is determined by the ontic state $\lambda$.
The third case can be treated in the same way and is discussed in the following subsection.
III.4 Separability and locality
Until now, we have introduced a general framework of an ontological theory with multiple realities. As previously said, the generalization to multiple instances is suggested by Bell’s theorem, which is proved under the assumption of single actual outcomes. Thus, we are particularly interested to formulate a separable and local framework. Here, we will consider the Bell scenario of two spacelike-separated laboratories (Case 3 of Sec. III.3). Each laboratory includes a system (${\cal Q}_{A}$ and ${\cal Q}_{B}$), a measuring device, and an observer – Alice and Bob. The device has a pointer that is set in a ready state and indicates an outcome $s$ after a measurement is performed on the system. The two systems ${\cal Q}_{A}$ and ${\cal Q}_{B}$ are generally in some entangled state. Eventually, Alice and Bob compare the results at a meeting point. Before the measurements, the instances are macroscopically indistinguishable, so that there is only one macroscopically distinct realization of Alice and Bob. As previously said, to avoid the creation of new instances over time, we assume that each spatially distinct object has the same number $m$ of instances.
Separability – Like in a separable one-world ontological theory, the two laboratories have their own ontic states, say
$\left\{\begin{array}[]{l}\lambda_{A}=(s_{A,1},s_{A,2},\dots,s_{A,m},\eta_{A}),\vskip 2.84526pt\\ \lambda_{B}=(s_{B,1},s_{B,2},\dots,s_{B,m},\eta_{B}).\end{array}\right$. (7)
It is important to emphasize again that the $m\times m$ possible combinations of local instances do not correspond to $m\times m$ instances of the overall system composed of the two laboratories. Each party has exactly $m$ instances at all times – including when they meet to compare their results. At the meeting point, one instance of Alice will interact with exactly one instance of Bob, and vice versa. Thus, given the one-to-one pairing rule, there will be $m$ pairs of Alice and Bob after they interact.
Locality – Initially, at time $t=0$, the ontic states are $\lambda_{A}^{\prime}=(s_{A,1}^{\prime},s_{A,2}^{\prime},\dots,s_{A,m}^{\prime},\eta_{A}^{\prime})$ and $\lambda_{B}^{\prime}=(s_{B,1}^{\prime},s_{B,2}^{\prime},\dots,s_{B,m}^{\prime},\eta_{B}^{\prime})$, $s_{A,k}^{\prime}$ and $s_{B,k}^{\prime}$ being in the ready state 0. Later, Alice and Bob perform each a measurement randomly selected among a set of possibilities. Let us denote by parameters $a$ and $b$ their respective selection. As usual in Bell’s scenario, we assume that $a$ and $b$ are not correlated with $\lambda_{A}^{\prime}$ and $\lambda_{B}^{\prime}$. These parameters are randomly generated outside the laboratories and passed to Alice and Bob. The ontic states evolve to $\lambda_{A}^{\prime\prime}=(s_{A,1}^{\prime\prime},s_{A,2}^{\prime\prime},\dots,s_{A,m}^{\prime\prime},\eta_{A}^{\prime\prime})$ and $\lambda_{B}^{\prime\prime}=(s_{B,1}^{\prime\prime},s_{B,2}^{\prime\prime},\dots,s_{B,m}^{\prime\prime},\eta_{B}^{\prime\prime})$, where $s_{A,k}^{\prime\prime}$ and $s_{B,k}^{\prime\prime}$ take the value of the outcome in the $k$-th instance of Alice and Bob, respectively.
At this point, we assume that the theory is locally causal bell76 . On Alice side we impose that the probability of getting $\lambda_{A}^{\prime\prime}$ given $a$, $b$, $\lambda_{A}^{\prime}$, $\lambda_{B}^{\prime}$ and $\lambda_{B}^{\prime\prime}$ does not depend on $b$, $\lambda_{B}^{\prime}$ and $\lambda_{B}^{\prime\prime}$. Similarly, we impose a specular condition on Bob side. That is,
$\begin{array}[]{l}\rho(\lambda_{A}^{\prime\prime}|\lambda_{B}^{\prime\prime},\lambda_{A}^{\prime},\lambda_{B}^{\prime},a,b)=\rho(\lambda_{A}^{\prime\prime}|\lambda_{A}^{\prime},a)\vskip 2.84526pt\\ \rho(\lambda_{B}^{\prime\prime}|\lambda_{A}^{\prime\prime},\lambda_{A}^{\prime},\lambda_{B}^{\prime},a,b)=\rho(\lambda_{B}^{\prime\prime}|\lambda_{B}^{\prime},b).\end{array}$ (8)
These conditions lead to the factorization
$\rho(\lambda_{A}^{\prime\prime},\lambda_{B}^{\prime\prime}|\lambda_{A}^{\prime},\lambda_{B}^{\prime},a,b)=\rho(\lambda_{A}^{\prime\prime}|\lambda_{A}^{\prime},a)\rho(\lambda_{B}^{\prime\prime}|\lambda_{B}^{\prime},b),$ (9)
from which Bell’s inequalities are derived in the single-world case bell76 . Generally, $\lambda_{A}^{\prime}$ and $\lambda_{B}^{\prime}$ are correlated, as the systems are entangled. Under the hypothesis of determinism, the assumption of local causality can be replaced by the weaker assumption of locality. See Ref. wiseman for a comprehensive discussion on the conceptual differences between the two assumptions in the context of Bell’s proofs.
Comparing the outcomes – After the measurements, Alice and Bob reach the meeting point and compare their results. The $i$-th instance of Alice is paired with the $k_{i}$-th instance of Bob, where $k_{i}=k_{i}(\lambda_{A}^{\prime\prime},\lambda_{B}^{\prime\prime})$ is some bijective function in $i$ depending on the ontic states $\lambda_{A}^{\prime\prime}$ and $\lambda_{B}^{\prime\prime}$. Depending on the model, the pairing may need only a partial information on $a$ and $b$, which is encoded in $\lambda_{A}^{\prime\prime}$ and $\lambda_{B}^{\prime\prime}$. Conversely, the DH theory needs the full information on the setting parameters (as shown later in Sec. IV). In Sec. V, we will see that the correlation of two maximally entangled qubits can be reproduced by a local model in which the two parties need to carry just one bit of information about their performed measurement to the meeting point. Furthermore, the model requires just two instances per party. In general, the number of realities and the amount of information can depend on the dimension of the Hilbert space of each subsystem, the number of subsystems and the computational network within which the subsystems interact.
In the next section, we will introduce a simplified version of the DH theory applied to the two-party Bell scenario, then we will introduce the $\psi$-epistemic model.
IV Nonlocal correlations in many-worlds theory
@@
In Ref. deutsch , Deutsch and Hayden proposed a local, many-worlds interpretation of quantum correlations by introducing a set of operators – referred to as descriptors – that evolve in the Heisenberg picture. This framework significantly expands the ontology of the theory by incorporating these new entities. Here, we present a highly simplified caricature of their argument. While this oversimplification may make the argument seem somewhat trivial, it nevertheless captures its essence. We consider Bell’s scenario with two local von Neumann measurements on two entangled qubits. Rather than using the descriptors, we describe each qubit by the operational parameters defining the measurements which are performed on them.
@@
The qubits are sent to two spatially separated parties, Alice and Bob. Each party independently chooses to perform a projective measurement, represented by the Bloch vectors $\vec{a}$ and $\vec{b}$, respectively. Alice gets $s_{A}=\pm 1$ and Bob $s_{B}=\pm 1$. The outcomes of each measurement occur with equal probability, that is,
$P(s_{A}|\vec{a})=P(s_{B}|\vec{b})=1/2.$ (10)
@@
In the DH framework, this means that the qubits locally split into two branches with equal weights. On each side, this process occurs without the need to know what is happening on the other side, as a consequence of the no signaling property of quantum theory. At this point of the experiment, no weight is attached on the pair of outcomes $(s_{A},s_{B})$. After the measurements, the parties decide to compare their results. To do that, they have to meet at the same place or send some communication through a classical channel. After multiple running of the experiment over different entangled states, they find out that the outcomes, $s_{A}\in\{-1,1\}$ and $s_{B}\in\{-1,1\}$, are distributed according to the probability distribution
$P(s_{A},s_{B}|\vec{a},\vec{b})=\frac{1}{4}\left(1-s_{A}s_{B}\vec{a}\cdot\vec{b}\right),$ (11)
which violate the Bell inequalities. While this violation leads to a break of local realism in a single-world scenario, the DH framework evades this conclusion as follows. When Alice meets with Bob, the branch in which she observed $s_{A}$ further divides into two branches labeled by the value $s_{B}$ observed by Bob. The probability of being paired with Bob observing $s_{B}$ given her observation $s_{A}$ is
$P(s_{B}|s_{A},\vec{a},\vec{b})\equiv P(s_{A},s_{B}|\vec{a},\vec{b})/P(s_{A}|\vec{a}).$ (12)
Thus, at the meeting point, four branches emerge such that each one is weighted with an amplitude whose modulus square is exactly $P(s_{A},s_{B}|\vec{a},\vec{b})$. By construction, these weights are not in conflict with the previous weight assignments. Rephrasing in Saunders’ terms, the (infinite) multitudes of microstates of each party are paired in such a way that the fraction of realities experiencing the outcomes $s_{A}$ and $s_{B}$ is equal to $P(s_{A},s_{B}|\vec{a},\vec{b})$. All these processes occur locally, which is guaranteed by the no signaling conditions
$\begin{array}[]{l}\sum_{s_{B}}P(s_{A},s_{B}|\vec{a},\vec{b})=P(s_{A}|\vec{a})\\ \sum_{s_{A}}P(s_{A},s_{B}|\vec{a},\vec{b})=P(s_{B}|\vec{b}).\end{array}$ (13)
It is worth recalling again Peres’ dictum that “unperformed measurements have no result”. The meeting is like a new measurement in which the pair of outcomes $(s_{A},s_{B})$ is made definite. As long as the comparison of the results does not occur, the parties can only claim their results (in their own reality), but the pairing of the two local realities remains indefinite. The idea of postponing the build-up of the correlations is illustrated in Ref. brassard for nonlocal PR boxes and further discussed in the general quantum case in Ref. charles . The model in Ref. brassard simplifies the pairing rule at the meeting point as the emerging branches have no weight. In particular, the rule is identical to the rule discussed in Sec. III.4: each instance of one party is paired with one instance of the other party.
Thus, the existence of many local realities allows us to delay the buildup of the correlations until the parties meet or communicate each other. It is fundamental to remark that, in the DH theory, each party has to carry the information about the performed measurement at the meeting point in order to get the right pairing. Namely, the probability $P(s_{A},s_{B}|\vec{a},\vec{b})$ must be known, which cannot be evaluated without knowing the Bloch vectors $\vec{a}$ and $\vec{b}$. Thus, each party has to communicate two real parameters, that is, infinite information. Employing the descriptors of Ref. deutsch , the number of real parameters would be $4^{2}-4=12$ per party charles2 . The total number of independent parameters is $15$, since the descriptors of Alice are not completely independent of the descriptors of Bob. This infinite amount of information is a consequence of the $\psi$-ontic nature of the theory.
V A model with finite information flow